4. You Corp, is analyzing two mutually exclusive projects. The free cash flows associated with these projects are as follows.10% DISCOUNT RATE
Answer #1

A.

Calculate the Projects payback period as follows:

2 Year Cash flows Cumulative cash flows 3 0 50.000.00 15,625.00 15,625.00 5 2 15.625.00 31,250.00 6 3 15.625.00 46,875.00 15,

AB 2 Year) Cash flows Cumulative cash flows 30 50000 4 1 15625 =B4 5 2 15625 =C4+B5 6 3 15625 =C5+B6 7 4 15625 =C6+B7 8 5 156

B & C.

Calculate the Net present value and internal rate of return as follows:

Year 0 Project 2 (50,000.00) B Project 1 (50,000.00 15.625.00 15.625.00 15.625.00 15,625.00 15.625.00 9,231.04 16.99% 8 Net p

20 1 Year Project 1 Project 2 -50000 -50000 15625 15625 5 3 15625 64 15625 15625 100000 8 Net present value (NPV) I=NPV(10%,B

D.

According to Payback period, Project 1 should be accepted as it has low payback period.

According to NPV, Project 2 should be accepted as it has higher NPV.

According to IRR, Project 1 should be accepted as it has higher IRR.